The mean hourly pay rate for financial managers in the East North Central region is 30$ and the variance is 9$^2 (Bureau of Labor Statistics). Assume that pay rates are normally distributed. 

a) Write the chance function of the above normal random variable.

b) What is the probability that a financial manager earns between 30$ and 35$ per hour? 

c) For a randomly selected financial manager, what is the probability the manager earned more than 20$ per hour?




1
Expert's answer
2021-08-26T17:16:01-0400

"N(30,9)"

a. Chance Function

"f(x)=\frac{1}{\sigma \sqrt{2\pi}}e^{-\frac{1}{2}(\frac{x-\mu}{\sigma})^2}, -\infin<x<\infin"

"f(x)=\frac{1}{3 \sqrt{2\pi}}e^{-\frac{1}{2}(\frac{x-30}{3})^2}, -\infin<x<\infin"

b. "P(30<x<35)"

"P(30<x<35)=P(\frac{30-30}{3}<Z<\frac{35-30}{3})"

"=P(0<Z<0.6)"

"=P(Z<0.6)-P(Z<0)"

"=0.7257-0.5"

"=0.2257"


c. "P(x>20)"

"P(x>20)=P(z>\frac{20-30}{3})"

"=1-P(Z<-3.333)"

"=1-0.004"

"=0.9996"